The vapor pressure of component A for a binary mixture (A and B) is given by the equation below:

PA = 6.1xAe-0.65xB + 1.6xB*xB

What is henry's law constant for component A?
1
Expert's answer
2020-12-18T08:41:18-0500

Henry's law:

C = k*P

k - henry's law constant

xA = k * (6.1xA-0.65xB + 1.6xB*xB)

xA = k * (6.1xA-0.65(1-xA) + 1.6*(1-xA)2)

"k=\frac{x_A}{6.1x_A-0.65(1-x_A) + 1.6*(1-x_A)^2}"